solve by addition. Indicate whether each system is independent, dependent or inconsistent.
-3x + 1y = 3
2x - 3y 5

Found 3 solutions by ewatrrr, stanbon, mananth:
Answer by ewatrrr(24785)


Hi
Solving by addition
-3x + 1y = 3 OR y = 3x+3
2x - 3y = 5
-9x + 3y = 3*3 |multiplying 1st equation thru by 3 to eliminate y by adding
2x - 3y = 5
-7x = 14
x = -2 and y = -3 [3*(-2) + 3]


Answer by stanbon(75887)

solve by addition. Indicate whether each system is independent, dependent or inconsistent.
-3x + 1y = 3
2x - 3y = 5
------
Multiply thru the 1st equation by 3:
-9x + 3y = 9
2x - 3y = 5
------
Add and solve for "x":
-7x = 14
x = -2
------
Solve for "y":
2x-3y = 5
2(-2)-3y = 5
-4 - 3y = 5
-3y = 9
y = -3
----------
The system has a single solution.
The system is independent.
===============================
Cheers,
Stan H.

Answer by mananth(16949)

-3x + 1y = 3
2x - 3y= 5
...
multiply first equation by 3
-9x+3y=9
add to second equation
-7x=14
x=-2
plug value of -2 in second equation
-4-3y=5
-3y=9
y=-3